# Numerical investigation of buoyancy-aided mixed convective flow past a square cylinder inclined at 45 degrees

### v3

@@ III.2 Time-Averaged Flow
Figure 15 reveals the variation of recirculation and inversion lengths with Ri. For calculation of the inversion length, time-averaged vorticity data is taken for all streamwise locations at a transverse location X $=0.1$. The location of the first sign change in the vorticity data yields the inversion length. To ensure independence of the result from the choice of transverse location, we have also calculated inversion length using X $=0.2$, and the results reveal no variation based on the choice of X. For the calculation of recirculation length, time-averaged streamwise velocity data along the centerline is considered. The location of the sign change in this data yields the recirculation length. There is a monotonic decrease in the inversion length up to Ri $=0.6$, followed by a marginal increase at Ri $=0.7$ and a subsequent reduction for higher Ri. The decrease in inversion length is due to the increasing strength of the natural convective vorticity over the strength of the forced convective vorticity with increasing Ri. The forced convective component is strong at low Ri but diminishes with downstream distance. At the same time, the natural convective component is weak in the near-field, but its strength grows with downstream distance. When Ri is increased, the strength of the natural convective vorticity closer to the cylinder also grows. The recirculation length shows an interesting variation, which was discussed briefly with Figure 13(a). It increases up to Ri $=0.7$, where it peaks, followed by a decrease. Prior to Ri $=0.7$, vortex shedding actively occurs in the flow. Due to the jet formed by buoyancy, the interaction between the two shear layers is inhibited, as seen in near-field plots of vorticity and temperature (see Figures 4 and 7). This increases the shear layer length and, consequently, the recirculation length, which increases up to Ri $=0.7$. Beyond Ri $=0.7$, there is no vortex shedding, and the near-field flow becomes steady. Since the flow is steady, there is no mutual interaction between the opposite shear layers, which form a steady separation bubble, and the effect of increasing buoyancy beyond the critical suppression value is to add momentum to the separation bubble, reducing its length.
@@
Figure 15: Variation of recirculation and inversion lengths with Richardson number. Recirculation length has been calculated by considering the time-averaged centerline (X $=0$) velocities, while the inversion length has been calculated by considering time-averaged velocities parallel to the centerline at X $=0.1$.
@@
Streamwise velocity, transverse velocity, vorticity, and temperature profiles at different streamwise locations for different Ri are shown in Figures 16, 17, 18 and 19, respectively. Figure 16 shows the deficit of momentum in the near-field associated with wake-like behavior and a recovery and subsequent excess of momentum in the intermediate and far-fields. In the near-field, the momentum deficit is seen for all Ri, resembling the profile seen in wakes. With increased downstream distance, the recovery of momentum occurs, starting earlier downstream for higher Ri, followed by lower Ri later downstream. At Y $=30$, the recovery of momentum is complete for all positive Ri, and there is an excess of momentum seen for higher Ri, which grows more pronounced further downstream, resembling the profile seen in jets. Therefore, the flow field downstream starts resembling a wake at low Ri and slowly transforms into that of a plume as Ri is increased, first far downstream, then near the cylinder. Such behavior is attributed to buoyancy, which tends to cause a jetting effect in the entire flow field, starting from the far downstream region and slowly moving towards the cylinder. Figure 17 shows the variation of transverse velocity, which is positive in the left half and negative in the right half of the flow for all positive Ri at all downstream locations. Figure 18 reveals the inversion of vorticity in the near and intermediate fields. For higher Ri, the inversion occurs earlier downstream, while for lower Ri, it occurs further downstream as the strength of the natural convective vorticity is lesser at lower Ri. The peak vorticity is seen to be higher for higher Ri. Figure 19 shows the temperature profiles for different streamwise locations. Overall, the temperature shows a dissipating trend with increased streamwise distance. We have also observed that prior to the onset distance of the far-field unsteadiness, the temperature profile width for the cases above the critical Ri is much lower than that of the cases below the critical Ri (see Figure 19(e)). However, after the onset distance of the far-field unsteadiness, the profile widens drastically, and it is comparable with the cases below the critical Ri (see Figure 19(f)).
(a) Y $=1$
(b) Y $=5$
(c) Y $=10$
(d) Y $=30$
(e) Y $=50$
(f) Y $=80$
Figure 16: Streamwise velocity profile at (a) Y $=1$, (b) Y $=5$, (c) Y $=10$, (d) Y $=30$, (e) Y $=50$, and (f) Y $=80$.
(a) Y $=1$
(b) Y $=5$
(c) Y $=10$
(d) Y $=30$
(e) Y $=50$
(f) Y $=80$
Figure 17: Transverse velocity profile at (a) Y $=1$, (b) Y $=5$, (c) Y $=10$, (d) Y $=30$, (e) Y $=50$, and (f) Y $=80$.
(a) Y $=1$
(b) Y $=5$
(c) Y $=10$
(d) Y $=30$
(e) Y $=50$
(f) Y $=80$
Figure 18: Vorticity profile at (a) Y $=1$, (b) Y $=5$, (c) Y $=10$, (d) Y $=30$, (e) Y $=50$, and (f) Y $=80$.
(a) Y $=1$
(b) Y $=5$
(c) Y $=10$
(d) Y $=30$
(e) Y $=50$
(f) Y $=80$
Figure 19: Temperature profile at (a) Y $=1$, (b) Y $=5$, (c) Y $=10$, (d) Y $=30$, (e) Y $=50$, and (f) Y $=80$.
@@ III.4 Unsteadiness of the Far-Field
Instantaneous and time-averaged contours in Figures 6, 3, 11 and 9 reveal that at Ri $=0.0$, the far-field is steady, while for Ri $>0.0$, the far-field is unsteady, which can be attributed to the increasing strength of the natural convective flow which makes the far-field behave like a plume. At the critical Ri and beyond, the far-field develops large-scale unsteadiness that is very similar to plume-like unsteadiness reported in the literature [25, 26, 27, 31]. This occurs in conjunction with the suppression of near-field vortex shedding. There is no range of Ri for which the flow field is entirely steady. As Ri is increased beyond the critical Ri, the point of onset of the far-field instability appears to come closer to the cylinder, which further indicates that the plume strength increases closer to the cylinder with increasing Ri. For Ri $\leq 0.6$, there is a widening of the plume in the intermediate-field itself, and in the far-field, the plume narrows (see Figures 11 and 9). For Ri $\geq 0.7$, the plume is narrow in the near and intermediate fields, and it sharply widens in the far-field with the onset of the far-field unsteadiness. The natural convective component grows stronger with increased downstream distance, and the opposite phenomenon occurs with the forced convective component. Due to this, there is a large momentum excess far downstream (Y $>$ 60), which causes strong jet-like behavior in the far-field. This can be seen in Figures 16 and 13(a). The jetting effect in the far-field is responsible for the plume-like unsteadiness of the far-field, which shows similar characteristics as the instabilities observed in natural convection plumes [25, 26, 27, 31]. The flow structures observed in the far-field require further detailed investigation to gain a complete understanding of flow physics.
The critical Ri for the transition to steady flow for both square and circular cylinders is 0.15, as reported by Sharma and Eswaran [21], and Chang and Sa [22]. However, the critical Ri is between 0.6 and 0.7 for our inclined square cylinder, which is significantly greater than that reported for square and circular cylinder cases. This large variation may be attributed to the (i) thickness and vorticity of the separation shear layer contained in it and/or (ii) afterbody effect, that is, the part of the obstacle beyond the separation of the flow from it. The shape of the obstacle and the corresponding afterbody are expected to affect the transition sequence in mixed convective flows past obstacles. The exact details of the effect require further investigation.
@@
[…]
@@
Figure 4 shows the total vorticity (flood) and the baroclinic vorticity (lines) in the near-field region. At Ri $=0.0$, there is no baroclinic vorticity due to the absence of thermal gradients. However, for positive Ri, baroclinic vorticity affects the near-field flow structures. The shed vortices appear to be stretched in the transverse direction, matching the shape of the baroclinic vortices. This effect is intensified at higher Ri. On comparing the shed vortices at Ri $=0.7$ with those at Ri $=0.0$, it is clear that the vortices at higher Ri are thinner and elongated compared to those at lower Ri. For Ri $<0.7$, the baroclinic vorticity follows an alternating pattern, with vortices of opposite signs being shed one after the other. The regions with a high magnitude of total vorticity are accompanied by positive baroclinic vorticity, while regions with a low magnitude of total vorticity are accompanied by negative baroclinic vorticity. This reveals that baroclinic vorticity increases the vorticity content in high-vorticity regions while cancelling the vorticity in the low-vorticity regions.
@@
Figure 5 shows the total vorticity (flood) and the baroclinic vorticity (lines) in the far-field region. It is seen that the baroclinic vorticity term plays a significant role in the development of the far-field unsteadiness. The baroclinic vorticity at Ri $=0.2$ creates unsteady flow structures in the far-field, unlike the forced convective flow case. It can also be seen that at this Ri, the effect of the baroclinic vorticity diminishes with increased downstream distance, as does the total vorticity. At Ri $=0.4$, the baroclinic vorticity is more pronounced. Regions of positive baroclinic vorticity increase the magnitude of total vorticity (in both positive and negative vorticity streams), while regions of negative baroclinic vorticity decrease the magnitude of total vorticity in both vorticity streams. A similar trend is observed for Ri $=0.6$. However, the strength of baroclinic vorticity increases for Ri beyond 0.7, which causes large-scale undulation of the shear layer, leading to unsteady flow structures. This effect is intensified at higher Ri. Similar to Ri $\leq$ 0.7, the regions of positive baroclinic vorticity enhance the magnitude of total vorticity, while regions of negative baroclinic vorticity minimize or cancel the magnitude of total vorticity. This behaviour is similar to that seen in the near-field.
@@
Figure 20 shows the streamwise variation of peak total vorticity and peak baroclinic vorticity for different Ri. For Ri $<0.7$, the total vorticity shows a decaying trend in the streamwise direction due to the diffusion of vorticity as it is convected away from the cylinder. Above Ri $=0.7$, the vorticity decays sharply after the cylinder and increases immediately thereafter, remaining constant until the onset of the far-field unsteadiness around Y $=65$. Both Ri of 0.7 and 1.0 show a maxima in the total vorticity with the onset of the far-field unsteadiness. This reveals that the shedding of vortices in the far-field occurs when the shear layer is unable to contain further vorticity beyond a prescribed limit. After the shedding begins, the vorticity once again decays with space due to the diffusion effect. On examining the plot of peak baroclinic vorticity, it is evident that a behavior similar to total vorticity is shown, with maxima in baroclinic vorticity visible in the far-field region.

### v4

@@ III.2 Time-Averaged Flow
Figure 15 reveals the variation of recirculation and inversion lengths with Ri. For the calculation of the inversion length, time-averaged vorticity data is taken for all streamwise locations at a transverse location X =0.1=0.1. The location of the first sign change in the vorticity data yields the inversion length. To ensure independence of the result from the choice of transverse location, we have also calculated inversion length using X =0.2=0.2, and the results reveal no variation based on the choice of X. For the calculation of recirculation length, time-averaged streamwise velocity data along the centerline are considered. The location of the sign change in this data yields the recirculation length. There is a monotonic decrease in the inversion length up to Ri =0.6=0.6, followed by a marginal increase at Ri =0.7=0.7 and a subsequent reduction for higher Ri. The decrease in inversion length is due to the increasing strength of the natural convective vorticity over the strength of the forced convective vorticity with increasing Ri. The forced convective component is strong at low Ri but diminishes with downstream distance. At the same time, the natural convective component is weak in the near-field, but its strength grows with downstream distance. When Ri is increased, the strength of the natural convective vorticity closer to the cylinder also grows. The recirculation length shows an interesting variation, which was discussed briefly with Figure 13(a). It increases up to Ri =0.7=0.7, where it peaks, followed by a decrease. Prior to Ri =0.7=0.7, vortex shedding actively occurs in the flow. Due to the jet formed by buoyancy, the interaction between the two shear layers is inhibited, as seen in near-field plots of vorticity and temperature (see Figures 4 and 7). This increases the shear layer length and, consequently, the recirculation length, which increases up to Ri =0.7=0.7. Beyond Ri =0.7=0.7, there is no vortex shedding, and the near-field flow becomes steady. Since the flow is steady, there is no mutual interaction between the opposite shear layers, which form a steady separation bubble, and the effect of increasing buoyancy beyond the critical suppression value is to add momentum to the separation bubble, reducing its length.
@@
Figure 15: Variation of recirculation and inversion lengths with Ri. Recirculation length has been calculated by considering the time-averaged centerline (X =0=0) velocities, while the inversion length has been calculated by considering time-averaged velocities parallel to the centerline at X =0.1=0.1.
@@
The streamwise velocity, transverse velocity, vorticity, and temperature profiles at different streamwise locations for different Ri are shown in Figures 16, 17, 18 and 19, respectively. Figure 16 shows the deficit of momentum in the near-field associated with wake-like behavior and a recovery and subsequent excess of momentum in the intermediate and far-fields. In the near-field, the momentum deficit is seen for all Ri, resembling the profile seen in wakes. With increased downstream distance, the recovery of momentum occurs, starting earlier downstream for higher Ri, followed by lower Ri later downstream. At Y =30=30, the recovery of momentum is complete for all positive Ri, and there is an excess of momentum seen for higher Ri, which grows more pronounced further downstream, resembling the profile seen in jets. Therefore, the flow field downstream starts resembling a wake at low Ri and slowly transforms into that of a plume as Ri is increased, first far downstream, then near the cylinder. Such behavior is attributed to buoyancy, which tends to cause a jetting effect in the entire flow field, starting from the far downstream region and slowly moving towards the cylinder. Figure 17 shows the variation of transverse velocity, which is positive in the left half and negative in the right half of the flow for all positive Ri at all downstream locations. Figure 18 reveals the inversion of vorticity in the near and intermediate fields. For higher Ri, the inversion occurs earlier downstream, while for lower Ri, it occurs further downstream as the strength of the natural convective vorticity is lesser at lower Ri. The peak vorticity is seen to be higher for higher Ri. Figure 19 shows the temperature profiles for different streamwise locations. Overall, the temperature shows a dissipating trend with increased streamwise distance. We have also observed that prior to the onset distance of the far-field unsteadiness, the temperature profile width for the cases above the critical Ri is much lower than that of the cases below the critical Ri (see Figure 19(e)). However, after the onset distance of the far-field unsteadiness, the profile widens drastically, and it is comparable with the cases below the critical Ri (see Figure 19(f)).
(a) Y =1=1
(b) Y =5=5
(c) Y =10=10
(d) Y =30=30
(e) Y =50=50
(f) Y =80=80
Figure 16: Streamwise velocity profile at (a) Y =1=1, (b) Y =5=5, (c) Y =10=10, (d) Y =30=30, (e) Y =50=50, and (f) Y =80=80.
(a) Y =1=1
(b) Y =5=5
(c) Y =10=10
(d) Y =30=30
(e) Y =50=50
(f) Y =80=80
Figure 17: Transverse velocity profile at (a) Y =1=1, (b) Y =5=5, (c) Y =10=10, (d) Y =30=30, (e) Y =50=50, and (f) Y =80=80.
(a) Y =1=1
(b) Y =5=5
(c) Y =10=10
(d) Y =30=30
(e) Y =50=50
(f) Y =80=80
Figure 18: Vorticity profile at (a) Y =1=1, (b) Y =5=5, (c) Y =10=10, (d) Y =30=30, (e) Y =50=50, and (f) Y =80=80.
(a) Y =1=1
(b) Y =5=5
(c) Y =10=10
(d) Y =30=30
(e) Y =50=50
(f) Y =80=80
Figure 19: Temperature profile at (a) Y =1=1, (b) Y =5=5, (c) Y =10=10, (d) Y =30=30, (e) Y =50=50, and (f) Y =80=80.
@@ III.4 Unsteadiness of the Far-Field
Instantaneous and time-averaged contours in Figures 6, 3, 11 and 9 reveal that at Ri =0.0=0.0, the far-field is steady, while for Ri >0.0>0.0, the far-field is unsteady, which can be attributed to the increasing strength of the natural convective flow which makes the far-field behave like a plume. At the critical Ri and beyond, the far-field develops large-scale unsteadiness that is very similar to plume-like unsteadiness reported in the literature [27, 28, 29, 42]. This occurs in conjunction with the suppression of near-field vortex shedding. There is no range of Ri for which the flow field is entirely steady. As Ri is increased beyond the critical Ri, the point of onset of the far-field instability appears to come closer to the cylinder, which further indicates that the plume strength increases closer to the cylinder with increasing Ri. For Ri $\leq 0.6$, there is a widening of the plume in the intermediate-field itself, and in the far-field, the plume narrows (see Figures 11 and 9). For Ri $\geq 0.7$, the plume is narrow in the near and intermediate fields, and it sharply widens in the far-field with the onset of the far-field unsteadiness. The natural convective component grows stronger with increased downstream distance, and the opposite phenomenon occurs with the forced convective component. Due to this, there is a large momentum excess far downstream (Y >> 60), which causes strong jet-like behavior in the far-field. This can be seen in Figures 16 and 13(a). The jetting effect in the far-field is responsible for the plume-like unsteadiness of the far-field, which shows similar characteristics as the instabilities observed in natural convection plumes [27, 28, 29, 42]. The flow structures observed in the far-field require further detailed investigation to gain a complete understanding of flow physics.
To add to the understanding of the transition from steady to unsteady flow in the far-field, we have analyzed the velocity profiles in the far-field and utilized Fjortoft’s theorem, discussed in detail by Arnal et al. [52], in order to understand the stability of the far-field flow in a clearer way. Fjortoft’s theorem utilizes points of inflexion in the time-averaged velocity profile to give a criterion for the stability of nearly parallel flows such as jets, wakes, and plumes, where the gradients of velocity in the transverse direction are significantly higher than those in the streamwise direction:
The point $X_{0}$ at which the velocity profile $V(X)$ satisfies the following condition is called the point of inflexion (PI):
$\left.\frac{\partial^{2}V}{\partial X^{2}}\right|_{X=X_{0}}=0$
According to Fjortoft’s theorem, a necessary (but not sufficient) condition for inviscid instability is that
$\frac{\partial^{2}V}{\partial X^{2}}(V-V_{0})<0$
must be true somewhere in the flow, where $V_{0}$ is the velocity at the inflexion point $X=X_{0}$. We have plotted the streamwise velocity profile against the transverse coordinate along with its first and second derivatives in Figure 20 at Ri =0.6=0.6. The transverse locations corresponding to $\overline{V}_{XX}=0$ intersect at X $\pm 2.75$, which are the two points of inflexion in the velocity profile, and the corresponding velocity at these points is $V_{0}$. It is clear from the graph that for $2.75\leq X\leq 2.75$, the quantity ($V-V_{0}$) is positive, while the quantity $\dfrac{\partial^{2}{V}}{\partial X^{2}}$ is negative. Consequently, one can say that the flow is unstable in the far field when subject to a suitable inviscid perturbation. The far field is stable at Ri =0.6=0.6, while it becomes unstable (unsteady) at Ri =0.7=0.7 because of the viscous flow in the present study.
Figure 20: Profiles of time-averaged streamwise velocity, first derivative, and second derivative at Y =80=80 for Ri =0.6=0.6. Points of inflexion in the velocity profile exist at X $\pm 2.75$.
The understanding of the transitional nature of the present mixed convective flow may become easier by dividing the flow downstream of the cylinder into two zones using the magnitude of the streamwise velocity along the centerline, as discussed below. Often, it has been discussed in the literature that flows of this nature are strictly forced convective below a certain Ri, and natural convective above a certain Ri. However, at each Ri, the flow in the present study is a spatially evolving one, similar to other transitional forced flows such as laminar-turbulent transitions in boundary layers, jets, etc., where there is no exact value of the control parameter(s) at which the transition of the entire flow occurs. In other words, one may expect, even at a particular Ri, that the flow has both forced and natural convective regimes in different zones of the flow. We have discussed earlier that the downstream flow near the cylinder is dominated by forced convection, while the flow far from the cylinder is dominated by natural convection. One may interpret the deficit in streamwise velocity along the centerline ($\overline{V}<1$) to be the forced convection-dominated zone (wake), while the excess in streamwise velocity ($\overline{V}>1$) may be understood as the natural convection-dominated zone (plume) for the present problem. For this purpose, we have plotted the wake length against Ri, where the wake length is defined as the downstream distance where the streamwise velocity along the centerline crosses unity (see Figure 21(a)). Therefore, the flow starts out as forced convective in the entire domain at Ri =0.0=$0.$ Subsequently, with increasing positive Ri, the wake length reduces, increasing the natural convection zone length. So, the transition between forced and natural convection is not only governed by Ri, but it is also a spatial phenomenon, as at each Ri, there is a region of the flow behaving like a wake, and the rest seems like a natural convective plume.
We have further analyzed the behaviour of the free convection-dominated far-field region. We have employed scaling analysis to show that the flow in the far-field is similar to that of a plume, as it shows self-similar behaviour. It is well known that plumes in the far-field of heat sources show self-similar behaviour in the time-averaged sense, similar to jets and wakes [53]. In order to scale the streamwise velocity, we have employed the local maximum velocity $V_{s}$, and to scale the transverse length X, we have employed the plume half-width ($L_{1/2}$), which is defined as the length measured from the centerline where the streamwise velocity is half of the local maximum velocity $V_{s}$. On plotting the scaled velocity profiles at different streamwise positions, it is found that they collapse onto a single dimensionless profile, proving a self-similarity behavior in the far-field region. The self-similar velocity follows a Gaussian distribution as given in Equation 10. The distribution is characterized by four constants whose values are listed in Table 9. The plot of the self-similar velocity along with the fitted Gaussian curve is provided in Figure 21(b) for Ri =0.7=0.7. The far field at a higher Ri =1.0=1.0 reveals similar self-similar behavior. Hunt and van den Bremer [53] also reported an identical Gaussian distribution.
$V=V_{0}+\dfrac{A}{w\sqrt{\pi/2}}e^{\dfrac{-2(X-X_{c})^{2}}{w^{2}}}$ (10)
The parameters of the equation that were obtained upon fitting the data of Ri =0.7=0.7 and 1.0 are given below in Table 9.
Figure 21: (a) Variation of wake length with Ri and (b) self-similar profile of streamwise velocity in the far-field for Ri =0.7=0.7. The remainder of the downstream domain beyond the wake length is the natural convection-dominated zone. The self-similar velocity profile is best described by a Gaussian curve, which has been fitted to the velocity data.
Table 9: Parameters of fitted curve for $Ri=0.7$ and $Ri=1.0$.
Parameter $Ri=0.7$ $Ri=1.0$
$V_{0}$ 0.431 0.388
$X_{c}$ -0.012 -0.015
$w$ 1.042 1.147
AA 0.752 0.892
@@
$\frac{\partial\omega}{\partial\tau}+\frac{\partial(\omega u_{i})}{\partial x_{j}}=\frac{1}{Re}\left[\frac{\partial^{2}\omega}{\partial x_{i}^{2}}\right]+Ri\frac{\partial\theta}{\partial x_{i}}\delta_{i2}$ (11)
@@
The baroclinic vorticity (lines) superimposed on the total vorticity (flood) in the near-field region has been depicted in Figure 4. At Ri =0.0=0.0, there is no baroclinic vorticity due to the absence of thermal gradients. However, for positive Ri, baroclinic vorticity affects the near-field flow structures. The shed vortices appear to be stretched in the transverse direction, matching the shape of the baroclinic vortices. This effect is intensified at higher Ri. On comparing the shed vortices at Ri =0.7=0.7 with those at Ri =0.0=0.0, it is clear that the vortices at higher Ri are thinner and elongated compared to those at lower Ri. For Ri <0.7<0.7, the baroclinic vorticity follows an alternating pattern, with vortices of opposite signs being shed one after the other. The regions with a high magnitude of total vorticity are accompanied by positive baroclinic vorticity, while regions with a low magnitude of total vorticity are accompanied by negative baroclinic vorticity. This reveals that baroclinic vorticity increases the vorticity content in high-vorticity regions while cancelling the vorticity in the low-vorticity regions.
@@
Figure 5 presents the baroclinic vorticity (lines) superimposed on the total vorticity (flood) in the far-field region. It is seen that the baroclinic vorticity term plays a significant role in the development of the far-field unsteadiness. The baroclinic vorticity at Ri =0.2=0.2 creates unsteady flow structures in the far-field, unlike the forced convective flow case. It can also be seen that at this Ri, the effect of the baroclinic vorticity diminishes with increased downstream distance, as does the total vorticity. At Ri =0.4=0.4, the baroclinic vorticity is more pronounced. Regions of positive baroclinic vorticity increase the magnitude of total vorticity (in both positive and negative vorticity streams), while regions of negative baroclinic vorticity decrease the magnitude of total vorticity in both vorticity streams. A similar trend is observed for Ri =0.6=0.6. However, the strength of baroclinic vorticity increases for Ri beyond 0.7, which causes large-scale undulation of the shear layer, leading to unsteady flow structures. This effect is intensified at higher Ri. Similar to Ri $\leq$ 0.7, the regions of positive baroclinic vorticity enhance the magnitude of total vorticity, while regions of negative baroclinic vorticity minimize or cancel the magnitude of total vorticity. This behaviour is similar to that seen in the near-field.
@@
Figure 22 shows the streamwise variation of peak total vorticity and peak baroclinic vorticity for different Ri. For Ri <0.7<0.7, the total vorticity shows a decaying trend in the streamwise direction due to the diffusion of vorticity as it is convected away from the cylinder. Above Ri =0.7=0.7, the vorticity decays sharply after the cylinder and increases immediately thereafter, remaining constant until the onset of the far-field unsteadiness around Y =65=65. Both Ri of 0.7 and 1.0 show a maxima in the total vorticity with the onset of the far-field unsteadiness. This reveals that the shedding of vortices in the far-field occurs when the shear layer is unable to contain further vorticity beyond a prescribed limit. After the shedding begins, the vorticity once again decays with space due to the diffusion effect. On examining the plot of peak baroclinic vorticity, it is evident that a behavior similar to total vorticity is shown, with maxima in baroclinic vorticity visible in the far-field region.